# A Variational Framework for Residual-Based Adaptivity in Neural PDE Solvers and Operator Learning

### v1

@@ 4.2 Physics-Informed Neural Networks (PINNs)
The total loss function combines the individual loss terms for the governing equations ($\mathcal{L}_{E}$), boundary, or initial conditions ($\mathcal{L}_{B}$), and, for inverse problems, observational data ($\mathcal{L}_{D}$):
@@ 4.3 Operator Learning
Let $\mathcal{X}$ be a space of functions over a domain $\Omega_{X}\subset\mathbb{R}^{d_{x}}$, and $\mathcal{Y}$ be a space of functions over $\Omega_{Y}\subset\mathbb{R}^{d_{y}}$. The operator of interest is:
@@
The goal is to learn a parametric model $G_{\theta}$ that approximates $\mathcal{G}$. The residual $R:\mathcal{X}\times\Omega_{y}\to\mathbb{R}^{+}$ for this task is defined as the difference between the operator’s prediction and the true solution:
@@
The loss function for a batch of $b_{u}$ functions is updated as follows:
@@
where the functions $\{v_{j}\}_{j=1}^{b_{u}}$ are sampled from the full set of training functions. The term $\Lambda$ is a matrix of importance weights, where $\Lambda_{i,j}$ corresponds to point $x_{i}$ for function $v_{j}$. These weights are constructed from a target p.m.f. matrix $Q^{k}$ constructed based on the choice of potential. For instance when $\Phi(x)=e^{x}$, $Q^{k}\in\mathbb{R}^{N\times N_{\text{func}}}$ is defined recursively as:
@@
Note that each column of the matrix $Q$ (for a fixed function $j$) is a p.m.f. over the spatial points, focusing attention on high-residual regions for that specific function. The weights are then smoothed over time with an EMA:
@@
A key advantage of this framework is that, if $\eta\neq 1-\gamma$, we can leverage the imbalance on learned spatial weights, $\Lambda_{i,j}$, to construct a sampling distribution over the functions themselves. The intuition is that functions with higher overall residuals will naturally accumulate larger $\Lambda$ values over time. Therefore, we propose the following approach to create a function-level sampling distribution. First, we compute an aggregated importance score $s_{j}$ for each function by summing its spatial weights:
@@ 5.1.1 Allen-Cahn Equation
Table 2: State-of-the-art (SOTA) comparison of relative $L^{2}$ errors for the Allen-Cahn and Burgers equations, focusing specifically on the performance of various quasi-Newton (second-order) optimization methods. The results from this work (TW) are benchmarked against reported results in recent literature. As described in the preceding sections, the Residual-based Adaptive Distribution (RAD) is a specific type of vRBA that utilizes a quadratic potential without smoothing. The results in this table, therefore, underscore that the combination of an adaptive sampling strategy with a suitable optimizer is critical for achieving high accuracy. Enhancements from the literature are abbreviated as follows: PirateNet (PN) [77], Fourier Features embeddings ($F$) [78], Weight Factorization (WF) [79], Causality (CS) [80], Learning Rate Annealing (LRA) [81] and Wolfe Line Search (WLS).
To place our results in context, we compare them against other state-of-the-art (SOTA) methods for the Allen-Cahn equation in Table 2. The comparison focuses on highly accurate solutions obtained with quasi-Newton optimizers. Our proposed method, combining the SSBroyden optimizer with vRBA, yields the most accurate result among the compared methods, achieving a relative $L^{2}$ error of $2.14\times 10^{-6}$. It is essential to note that this performance is achieved with a minimal set of enhancements, only Fourier Features and our adaptive sampling framework. This contrasts with other approaches that rely on a more extensive suite of techniques yet yield less accurate results. Furthermore, the table highlights that the previous best result for this problem also uses an adaptive method, RAD. As detailed in our theoretical framework, RAD is a specific instance of vRBA that uses a quadratic potential without smoothing. This underscores our main conclusion: the combination of a powerful optimizer with a principled adaptive sampling strategy like vRBA is the critical factor for achieving the highest accuracy.
@@ 5.1.2 Burgers Equation
The Burgers’ equation is defined as:
@@
where $u$ represents the velocity field, subject to the dynamic viscosity $\nu=1/(100\pi)$. The initial condition and boundary conditions are described as follows:
@@ 5.2.1 Bubble Growth Dynamics-DeepONet
We study the dynamics of a single gas bubble in an incompressible liquid governed by the Rayleigh–Plesset (R–P) equation [82], a nonlinear ordinary differential equation describing the evolution of the bubble radius $R(t)$ under a time-varying pressure field $P_{\infty}(t)$. Under isothermal assumptions and negligible temperature variations, the simplified linearized R–P equation is:
@@ Acknowledgements
We acknowledge the support of the NIH grant R01AT012312, MURI/AFOSR FA9550-20-1-0358 project, the DOE-MMICS SEA-CROGS DE-SC0023191 award, and the ONR Vannevar Bush Faculty Fellowship (N00014-22-1-2795). We are also deeply grateful to Prof. Jerome Darbon for his insightful guidance on variational methods and optimization.
@@ Appendix A Representation of divergences
The first is the Laplace principle, which rewrites the maximum as integration against increasingly singular measures. While the result is well-known in the context of large deviations, e.g., [64, Theorem 4.3.1] or [66, Theorem 1.5], the statement as presented is difficult to find, so we provide the proof for completeness.
@@
A proof can be found in, for example, [66, Proposition 2.2]. It consists of two steps: first, showing that the identified $q^{*}$ achieves equality. Then, decompose the log-likelihood-ratio into ones with respect to the optimizer $q^{*}$ allows one to use the non-negativity of relative entropy to conclude.
In fact, the Gibbs variational principle is only one part of a duality with the log-partition function. This is known as the Donsker-Varadhan representation, e.g., [66, Lemma 2.4], which gives a representation of the relative entropy as the Legendre-Fenchel transform of the log-partition function:
@@
The first item follows from [71, Corollary 58]. The proof is involved and will not be reproduced here.
@@ D.3 JAX Implementation of the SSBroyden Optimizer
Our implementation preserves the core SSBroyden update logic, which dynamically computes scaling ($\tau_{k}$) and updating ($\phi_{k}$) parameters. However, the line search portion of the algorithm required a complete rewrite. Due to the absence of SciPy’s advanced line search routines in JAX, we developed a custom three-stage fallback line search mechanism to ensure robust convergence. This procedure creates a cascade of attempts with progressively less strict Wolfe conditions, starting with strict parameters ($c_{2}=0.9$) and relaxing them ($c_{2}=0.8$, then $c_{2}=0.5$) only upon failure. This adaptation was critical for ensuring the optimizer could consistently make progress on the challenging loss landscapes of the problems studied.

### v2

@@ 4.2 Physics-Informed Neural Networks (PINNs)
The total loss function combines the individual loss terms for the governing equations ($\mathcal{L}_{E}$), boundary, or initial conditions ($\mathcal{L}_{B}$), and, for inverse problems, observational data ($\mathcal{L}_{D}$), which reads
@@ 4.3 Operator Learning
Let $\mathcal{X}$ be a space of functions over a domain $\Omega_{X}\subset\mathbb{R}^{d_{x}}$, and $\mathcal{Y}$ be a space of functions over $\Omega_{Y}\subset\mathbb{R}^{d_{y}}$. The operator of interest is
@@
The goal is to learn a parametric model $G_{\theta}$ that approximates $\mathcal{G}$. The residual $R:\mathcal{X}\times\Omega_{y}\to\mathbb{R}^{+}$ for this task is defined as the difference between the operator’s prediction and the true solution and reads
@@
The loss function for a batch of $b_{u}$ functions is updated as follows
@@
where the functions $\{v_{j}\}_{j=1}^{b_{u}}$ are sampled from the full set of training functions. The term $\Lambda$ is a matrix of importance weights, where $\Lambda_{i,j}$ corresponds to point $x_{i}$ for function $v_{j}$. These weights are constructed from a target p.m.f. matrix $Q^{k}$ constructed based on the choice of potential. For instance when $\Phi(x)=e^{x}$, $Q^{k}\in\mathbb{R}^{N\times N_{\text{func}}}$ is defined recursively as follows
@@
Note that each column of the matrix $Q$ (for a fixed function $j$) is a p.m.f. over the spatial points, focusing attention on high-residual regions for that specific function. The weights are then smoothed over time with an EMA
@@
A key advantage of this framework is that, if $\eta\neq 1-\gamma$, we can leverage the imbalance on learned spatial weights, $\Lambda_{i,j}$, to construct a sampling distribution over the functions themselves. The intuition is that functions with higher overall residuals will naturally accumulate larger $\Lambda$ values over time. Therefore, we propose the following approach to create a function-level sampling distribution. First, we compute an aggregated importance score $s_{j}$ for each function by summing its spatial weights
@@ 5.1.1 Allen-Cahn Equation
Table 2: State-of-the-art comparison of relative $L^{2}$ errors for the Allen-Cahn and Burgers equations, focusing specifically on the performance of various quasi-Newton (second-order) optimization methods. The results from this work (TW) are benchmarked against reported results in recent literature. As described in the preceding sections, the Residual-based Adaptive Distribution (RAD) is a specific type of vRBA that utilizes a quadratic potential without smoothing. The results in this table, therefore, underscore that the combination of an adaptive sampling strategy with a suitable optimizer is critical for achieving high accuracy. Enhancements from the literature are abbreviated as follows: PirateNet (PN) [77], Fourier Features embeddings ($F$) [78], Weight Factorization (WF) [79], Causality (CS) [80], Learning Rate Annealing (LRA) [81] and Wolfe Line Search (WLS).
To place our results in context, we compare them against other state-of-the-art methods for the Allen-Cahn equation in Table 2. The comparison focuses on highly accurate solutions obtained with quasi-Newton optimizers. Our proposed method, combining the SSBroyden optimizer with vRBA, yields the most accurate result among the compared methods, achieving a relative $L^{2}$ error of $2.14\times 10^{-6}$. It is essential to note that this performance is achieved with a minimal set of enhancements, only Fourier Features and our adaptive sampling framework. This contrasts with other approaches that rely on a more extensive suite of techniques yet yield less accurate results. Furthermore, the table highlights that the previous best result for this problem also uses an adaptive method, RAD. As detailed in our theoretical framework, RAD is a specific instance of vRBA that uses a quadratic potential without smoothing. This underscores our main conclusion: the combination of a powerful optimizer with a principled adaptive sampling strategy like vRBA is the critical factor for achieving the highest accuracy.
@@ 5.1.2 Burgers Equation
The Burgers’ equation is defined as
@@
where $u$ represents the velocity field, subject to the dynamic viscosity $\nu=1/(100\pi)$. The initial condition and boundary conditions are described as follows
@@ 5.2.1 Bubble Growth Dynamics-DeepONet
We study the dynamics of a single gas bubble in an incompressible liquid governed by the Rayleigh–Plesset (R–P) equation [82], a nonlinear ordinary differential equation describing the evolution of the bubble radius $R(t)$ under a time-varying pressure field $P_{\infty}(t)$. Under isothermal assumptions and negligible temperature variations, the simplified linearized R–P equation reads
@@ Acknowledgements
We acknowledge the support of the NIH grant R01AT012312, MURI/AFOSR FA9550-20-1-0358 project, the DOE-MMICS SEA-CROGS DE-SC0023191 award, and the ONR Vannevar Bush Faculty Fellowship (N00014-22-1-2795).
@@ Appendix A Representation of divergences
The first is the Laplace principle, which rewrites the maximum as integration against increasingly singular measures. While the result is well-known in the context of large deviations, e.g., [65, Theorem 4.3.1] or [67, Theorem 1.5], the statement as presented is difficult to find, so we provide the proof for completeness.
@@
A proof can be found in, for example, [67, Proposition 2.2]. It consists of two steps: first, showing that the identified $q^{*}$ achieves equality. Then, decompose the log-likelihood-ratio into ones with respect to the optimizer $q^{*}$ allows one to use the non-negativity of relative entropy to conclude.
In fact, the Gibbs variational principle is only one part of a duality with the log-partition function. This is known as the Donsker-Varadhan representation, e.g., [67, Lemma 2.4], which gives a representation of the relative entropy as the Legendre-Fenchel transform of the log-partition function:
@@
The first item follows from [72, Corollary 58]. The proof is involved and will not be reproduced here.
@@ D.3 JAX Implementation of the SSBroyden Optimizer
Our implementation preserves the core SSBroyden update logic, which dynamically computes scaling ($\tau_{k}$) and updating ($\phi_{k}$) parameters. However, the line search portion of the algorithm required a complete rewrite. Due to the absence of SciPy’s advanced line search routines in JAX Scipy, we developed a custom three-stage fallback line search mechanism to ensure robust convergence. This procedure creates a cascade of attempts with progressively less strict Wolfe conditions, starting with strict parameters ($c_{2}=0.9$) and relaxing them ($c_{2}=0.8$, then $c_{2}=0.5$) only upon failure. This adaptation was critical for ensuring the optimizer could consistently make progress on the challenging loss landscapes of the problems studied.